OpenStudy (anonymous):

marlene rides her bicycle to her friend jons house and returns home by the same route. shes rides her bike at a constant speed on 6mph on level ground, 4mph while going uphill and 12mph going downhill. if her total time riding was 1 hr, how far is it to jons house?

OpenStudy (anonymous):

3 Miles r t = d, t = d/r Let L, a and b be the length of level ground, uphill length and downhill length to jon's house respectively. The following expression represents the total time for a round trip to jon's house:\[\frac{L}{6}+\frac{a}{4}+\frac{b}{12}+\frac{L}{6}+\frac{a}{12}+\frac{b}{4}=1\]Combine the factions in the above expression and simplify .\[a+b+L= 3\]